# On Selecting Few-Shot Examples for LLM-based Code Vulnerability Detection

### v1

@@ 5. Experiments
  1. (1)
    How do the proposed algorithms compare individually with baselines (zero shot and few shot settings) in helping large language models find vulnerabilities in code?
  2. (2)
@@
For datasets, we consider established benchmarks such as PrimeVul, DiverseVul, SVEN (Ding et al., 2024; Chen et al., 2023; He and Vechev, 2023). These are well-curated datasets, including pairs of code samples (vulnerable vs. non-vulnerable). We experiment with adding both vulnerable and non-vulnerable few-shot examples to better gauge the performance of the LLMs on the vulnerability detection task. We also leverage recent work on vulnerabilities in JavaScript programs (Cassel et al., 2023, 2025) and obtained a copy of the dataset generated by their tools directly from the authors. We refer to this dataset as NodeMedic.
@@
DiverseVul comprises 330492 unpaired samples, and it has 150 unique CWEs. To ensure computational tractability, the dataset was first partitioned into a primary training and a primary test set, following an 80:20 ratio which were unbalanced. Subsequently, based on the intersection of CWEs present in both splits, a final balanced sample was created. This resulted in a training set of 200 examples and a test set of 300 examples. Both of these sampled sets have 114 unique and comon CWEs.
SVEN comprises 1440 training and 166 validation samples. This dataset was partitioned into two distinct subsets based on the programming language of the functions. The C/C++ subset, designated SVENC, consists of 756 training and 90 validation samples. The Python subset, SVENP, is composed of 684 training and 76 validation samples. SVENC and SVENP has 7 and 4 unique CWEs respectively (both train and validation set).
NodeMedic dataset was provided by the authors of NodeMedic-FINE (Cassel et al., 2025), a dynamic analysis tool that detects taint flows from package APIs to dangerous sinks that may enable arbitrary command injection or code execution. The dataset is divided into 1,506 training and 189 test samples, each corresponding to a Node.js package with potentially vulnerable dataflows reported by the tool. All reports are either automatically confirmed by NodeMedic-FINE (Cassel et al., 2025) or manually verified by its authors.
To facilitate semantic code retrieval, we employed a specialized encoder model from the Salesforce SFR family: ”Salesforce/SFR-Embedding-Code-400M_R” (Liu et al., 2024). These vector embeddings are used in the nearest neighbor search, which utilizes cosine distance to identify the closest matches for any given query.
We conducted experiments on two open-source models, ”Qwen-2.5-Coder-7B-Instruct” (Hui et al., 2024) and ”gemma-3-4b-it” (Team, 2025), as well as the closed-source GPT-5-mini (the gpt-5-mini-2025-08-07 snapshot), to assess the generalizability of our proposed techniques. The Qwen model was selected for its strong, well-documented proficiency on coding tasks and very large context window, while the Gemma model was chosen for its computational efficiency and competitive performance given its parameter count. GPT-5-mini serves as a high-quality closed-source baseline to contextualize the behavior of the open models. The combination of these models allows evaluation across differing trade-offs of capability and resource requirements, while staying within our available computational budget.
Table 1. Results for all approaches (ZS = zero-shot, R-FS = random few-shot, LFM = Learn-from-Mistakes, LFNN = Learn-from-Nearest-Neighbors, CM = Combined Method) across three models. Gemma and Qwen results are mean over five runs; GPT results are based on one run. A dash (‘–’) indicates that the metric is undefined in at least one of the runs due to division by zero.
@@ Results
We present a detailed analysis of the experimental results from our evaluation of three LLMs from the Gemma, GPT, and Qwen families across the four datasets (Results are presented separately for SVENC and SVENP). All findings are summarized in Table 1, which reports the mean performance metrics over five independent runs for the Gemma and Qwen models using five different random seeds. The standard deviations between runs are minimal, so they are not included in the table because of space constraints. For the GPT model, we report the results from a single run. Our analysis is structured on a per-dataset basis to highlight the varying effectiveness of each few-shot selection approach under different data distributions and model capabilities.
@@
Finally, one interesting insight was hard examples (LFM) in context try to bias the model towards greater recall and lower precision. As CM3 applies LFM on LFNN examples keeping first LFM’s result as context, it also keep that trend mostly.
Quality of Vulnerability Detection Datasets
Ding et al. (2024) recently highlighted key challenges in existing vulnerability detection datasets, including label noise, data duplication, and data leakage. To mitigate these concerns, we incorporate the PrimeVul dataset in our experiments, as it was carefully curated to address such issues, along with several other commonly used datasets. We also include an additional unpublished dataset (NodeMedic) to demonstrate the generality of our approach. Beyond the noisiness of current vulnerability datasets, we acknowledge that function-level vulnerability detection has inherent limitations compared to repository-level detection (Risse et al., 2025). However, we believe that function-level datasets and detectors provide a valuable first step toward addressing vulnerability detection in broader contexts.
6. Conclusion
Appendix A Prompt Used for Vulnerability Detection
The prompt shown here is used when running LFM as well as for evaluating the vulnerability detection capabilities of the LLM for the experiments reported in Section 5. To instruct the model clearly, we provided a concise and explicit system instruction, guiding the model to behave strictly as a security expert and to output responses in a standardized format. The exact system-level instruction used to prime the LLM is defined as follows:
Appendix B Full Results with Standard Deviation
Table 2. Full results with standard deviation for all approaches across three models. Each cell reports the mean and standard deviation over five runs for Gemma and Qwen models, while GPT results are based on a single run and do not have standard deviation. For the approach names, the following abbreviations are used: ZS = zero-shot, R-FS = random few-shot, LFM = Learn-from-Mistakes, LFNN = Learn-from-Nearest-Neighbors, CM = Combined Method. A dash (‘–’) indicates that the metric is undefined in at least one of the runs due to division by zero. Dataset Approach Gemma-3-4b-it GPT-5-mini Qwen-2.5-Coder Acc Prec Recall F1 Acc Prec Recall F1 Acc Prec Recall F1 DiverseVul ZS 0.619 (±0.003) 0.770 (±0.009) 0.340 (±0.000) 0.472 (±0.002) 0.593 0.588 0.627 0.607 0.497 (±0.000) 0.000 (±0.000) 0.000 (±0.000) 0.000 (±0.000) R-FS 0.545 (±0.053) 0.571 (±0.081) 0.467 (±0.210) 0.480 (±0.132) 0.560 0.598 0.367 0.455 0.599 (±0.045) 0.741 (±0.027) 0.303 (±0.139) 0.411 (±0.150) LFM 0.525 (±0.003) 0.840 (±0.056) 0.061 (±0.005) 0.114 (±0.009) 0.617 0.647 0.513 0.573 0.500 (±0.000) 0.500 (±0.000) 1.000 (±0.000) 0.667 (±0.000) LFNN 0.512 (±0.004) 0.511 (±0.004) 0.537 (±0.009) 0.524 (±0.006) 0.590 0.624 0.453 0.525 0.659 (±0.009) 0.632 (±0.007) 0.759 (±0.014) 0.690 (±0.009) CM 1 0.548 (±0.020) 0.530 (±0.015) 0.868 (±0.035) 0.658 (±0.007) 0.587 0.610 0.480 0.537 0.638 (±0.008) 0.606 (±0.011) 0.792 (±0.039) 0.686 (±0.011) CM 2 0.515 (±0.002) 0.508 (±0.001) 0.973 (±0.000) 0.667 (±0.001) 0.560 0.573 0.473 0.518 0.597 (±0.006) 0.557 (±0.003) 0.949 (±0.008) 0.702 (±0.004) CM 3 0.499 (±0.002) 0.499 (±0.001) 0.985 (±0.007) 0.663 (±0.002) 0.570 0.585 0.480 0.527 0.531 (±0.021) 0.517 (±0.012) 0.988 (±0.015) 0.678 (±0.007) NodeMedic ZS 0.506 (±0.003) 0.765 (±0.005) 0.460 (±0.000) 0.574 (±0.001) 0.757 0.786 0.912 0.845 0.375 (±0.009) 0.870 (±0.049) 0.162 (±0.005) 0.273 (±0.008) R-FS 0.687 (±0.060) 0.726 (±0.003) 0.914 (±0.136) 0.804 (±0.060) 0.794 0.845 0.876 0.860 0.632 (±0.060) 0.748 (±0.020) 0.750 (±0.163) 0.739 (±0.071) LFM 0.725 (±0.000) 0.725 (±0.000) 1.000 (±0.000) 0.840 (±0.000) 0.767 0.789 0.927 0.852 0.720 (±0.000) 0.723 (±0.000) 0.993 (±0.000) 0.837 (±0.000) LFNN 0.758 (±0.004) 0.759 (±0.003) 0.975 (±0.004) 0.854 (±0.002) 0.788 0.849 0.861 0.855 0.713 (±0.011) 0.768 (±0.004) 0.866 (±0.014) 0.814 (±0.008) CM 1 0.701 (±0.005) 0.755 (±0.009) 0.870 (±0.018) 0.808 (±0.004) 0.788 0.839 0.876 0.857 0.751 (±0.008) 0.776 (±0.004) 0.923 (±0.020) 0.843 (±0.007) CM 2 0.565 (±0.008) 0.719 (±0.005) 0.657 (±0.008) 0.686 (±0.006) 0.794 0.840 0.883 0.861 0.735 (±0.006) 0.748 (±0.004) 0.958 (±0.003) 0.840 (±0.003) CM 3 0.722 (±0.007) 0.728 (±0.003) 0.984 (±0.014) 0.837 (±0.005) 0.794 0.846 0.883 0.864 0.720 (±0.000) 0.724 (±0.001) 0.991 (±0.003) 0.837 (±0.000) PrimeVul ZS 0.599 (±0.044) 0.723 (±0.106) 0.358 (±0.036) 0.472 (±0.003) 0.535 0.524 0.750 0.617 0.500 (±0.000) 0.002 (±0.004) 0.004 (±0.008) R-FS 0.522 (±0.041) 0.553 (±0.086) 0.373 (±0.148) 0.422 (±0.101) 0.560 0.556 0.600 0.577 0.503 (±0.010) 0.528 (±0.058) 0.232 (±0.105) 0.304 (±0.094) LFM 0.516 (±0.013) 0.726 (±0.186) 0.437 (±0.459) 0.336 (±0.270) 0.550 0.557 0.490 0.521 0.500 (±0.000) 0.500 (±0.000) 1.000 (±0.000) 0.667 (±0.000) LFNN 0.527 (±0.030) 0.520 (±0.017) 0.648 (±0.147) 0.571 (±0.064) 0.555 0.549 0.620 0.582 0.483 (±0.009) 0.390 (±0.056) 0.060 (±0.009) 0.104 (±0.015) CM 1 0.543 (±0.032) 0.527 (±0.022) 0.885 (±0.040) 0.659 (±0.010) 0.575 0.577 0.600 0.588 0.473 (±0.007) 0.462 (±0.010) 0.336 (±0.092) 0.383 (±0.057) CM 2 0.516 (±0.001) 0.508 (±0.001) 0.964 (±0.022) 0.666 (±0.005) 0.555 0.549 0.610 0.578 0.455 (±0.012) 0.445 (±0.015) 0.370 (±0.026) 0.404 (±0.021) CM 3 0.500 (±0.002) 0.500 (±0.001) 0.982 (±0.017) 0.663 (±0.004) 0.555 0.546 0.650 0.594 0.499 (±0.006) 0.499 (±0.003) 0.986 (±0.015) 0.663 (±0.005) SVENC ZS 0.478 (±0.000) 0.474 (±0.000) 0.400 (±0.000) 0.434 (±0.000) 0.533 0.523 0.756 0.618 0.500 (±0.000) 0.000 (±0.000) 0.000 (±0.000) R-FS 0.506 (±0.010) 0.503 (±0.005) 0.778 (±0.347) 0.564 (±0.174) 0.600 0.579 0.733 0.647 0.491 (±0.013) 0.422 (±0.373) 0.345 (±0.274) LFM 0.500 (±0.000) 0.500 (±0.000) 1.000 (±0.000) 0.667 (±0.000) 0.544 0.543 0.556 0.549 0.500 (±0.000) 0.500 (±0.000) 1.000 (±0.000) 0.667 (±0.000) LFNN 0.471 (±0.009) 0.428 (±0.034) 0.187 (±0.044) 0.258 (±0.049) 0.567 0.554 0.689 0.614 0.504 (±0.009) 0.567 (±0.133) 0.040 (±0.009) 0.075 (±0.017) CM 1 0.520 (±0.008) 0.513 (±0.005) 0.813 (±0.027) 0.629 (±0.010) 0.578 0.566 0.667 0.612 0.520 (±0.019) 0.527 (±0.024) 0.387 (±0.074) 0.442 (±0.055) CM 2 0.536 (±0.004) 0.520 (±0.003) 0.947 (±0.011) 0.671 (±0.002) 0.589 0.574 0.689 0.626 0.493 (±0.005) 0.495 (±0.004) 0.636 (±0.011) 0.556 (±0.005) CM 3 0.500 (±0.000) 0.500 (±0.000) 0.938 (±0.029) 0.652 (±0.007) 0.611 0.596 0.689 0.639 0.496 (±0.005) 0.498 (±0.003) 0.982 (±0.009) 0.661 (±0.004) SVENP ZS 0.587 (±0.006) 0.597 (±0.007) 0.537 (±0.013) 0.565 (±0.009) 0.763 0.727 0.842 0.780 0.705 (±0.011) 0.965 (±0.028) 0.426 (±0.020) 0.591 (±0.019) R-FS 0.558 (±0.038) 0.561 (±0.044) 0.632 (±0.117) 0.584 (±0.035) 0.789 0.806 0.763 0.784 0.616 (±0.027) 0.707 (±0.071) 0.432 (±0.144) 0.514 (±0.102) LFM 0.500 (±0.000) 0.500 (±0.000) 0.905 (±0.013) 0.644 (±0.003) 0.776 0.744 0.842 0.790 0.500 (±0.000) 0.500 (±0.000) 1.000 (±0.000) 0.667 (±0.000) LFNN 0.634 (±0.005) 0.641 (±0.004) 0.611 (±0.011) 0.625 (±0.007) 0.895 0.941 0.842 0.889 0.792 (±0.005) 0.845 (±0.002) 0.716 (±0.011) 0.775 (±0.007) CM 1 0.553 (±0.012) 0.546 (±0.009) 0.616 (±0.043) 0.579 (±0.023) 0.882 0.892 0.868 0.880 0.787 (±0.023) 0.793 (±0.034) 0.779 (±0.021) 0.785 (±0.019) CM 2 0.632 (±0.000) 0.639 (±0.000) 0.605 (±0.000) 0.622 (±0.000) 0.895 0.917 0.868 0.892 0.697 (±0.019) 0.660 (±0.016) 0.816 (±0.029) 0.729 (±0.018) CM 3 0.521 (±0.013) 0.512 (±0.008) 0.942 (±0.045) 0.663 (±0.009) 0.803 0.795 0.816 0.805 0.600 (±0.018) 0.558 (±0.012) 0.963 (±0.027) 0.707 (±0.010)
An extended version of Table 1 with standard deviation is shown in Table 2. A key observation is the general consistency of the methods—most exhibit low standard deviations. Table  2 reinforces the reliability of the mean results discussed in the main text.
Appendix C Different Variants of Learn-from-Mistakes (LFM)
@@
Table 3 shows a comprehensive breakdown of performance across different LFM configurations. It is evident that the method’s effectiveness is highly sensitive to its parameterization and the specific characteristics of the dataset.

### v2

@@ V Experiments
  1. 1.
    How do the proposed algorithms compare individually with baselines (zero-shot and few-shot settings) in helping large language models find vulnerabilities in code?
  2. 2.
@@
For datasets, we consider established benchmarks such as PrimeVul, DiverseVul, SVEN [8, 63, 64]. These are well-curated datasets, including pairs of code samples (vulnerable vs. non-vulnerable). We experiment with adding both vulnerable and non-vulnerable few-shot examples to better gauge the performance of the LLMs on the vulnerability detection task. We also leverage recent work on vulnerabilities in JavaScript programs [16, 17] and obtained a copy of the dataset generated by their tools directly from the authors. We refer to this dataset as NodeMedic.
@@
DiverseVul comprises 330492 unpaired samples, and it has 150 unique CWEs. To ensure computational tractability, the dataset was first partitioned into a primary training and a primary test set, following an 80:20 ratio which were unbalanced. Subsequently, based on the intersection of CWEs present in both splits, a final balanced sample was created. This resulted in a training set of 200 examples and a test set of 300 examples. Both of these sampled sets have 114 unique and common CWEs.
SVEN comprises 1440 training and 166 validation samples. This dataset was partitioned into two distinct subsets based on the programming language of the functions. The C/C++ subset, designated SVENC, consists of 756 training and 90 validation samples. The Python subset, SVENP, is composed of 684 training and 76 validation samples. SVENC and SVENP have 7 and 4 unique CWEs, respectively (both train and validation set).
NodeMedic dataset was provided by the authors of NodeMedic-FINE [17], a dynamic analysis tool that detects taint flows from package APIs to dangerous sinks that may enable arbitrary command injection or code execution. The dataset is divided into 1,506 training and 189 test samples, each corresponding to a Node.js package with potentially vulnerable dataflows reported by the tool. All reports are either automatically confirmed by NodeMedic-FINE [17] or manually verified by its authors.
To facilitate semantic code retrieval, we employed a specialized encoder model from the Salesforce SFR family: ”Salesforce/SFR-Embedding-Code-400M_R” [65]. These vector embeddings are used in the nearest neighbor search, which utilizes cosine distance to identify the closest matches for any given query.
We conducted experiments on two open-source models, ”Qwen-2.5-Coder-7B-Instruct” [66] and ”gemma-3-4b-it” [67], as well as the closed-source GPT-5-mini (the gpt-5-mini-2025-08-07 snapshot), to assess the generalizability of our proposed techniques. The Qwen model was selected for its strong, well-documented proficiency on coding tasks and very large context window, while the Gemma model was chosen for its computational efficiency and competitive performance given its parameter count. GPT-5-mini serves as a high-quality closed-source baseline to contextualize the behavior of the open models. The combination of these models allows evaluation across differing trade-offs of capability and resource requirements, while staying within our available computational budget.
TABLE I: Results for all approaches (ZS = zero-shot, R-FS = random few-shot, LFM = Learn-from-Mistakes, LFNN = Learn-from-Nearest-Neighbors, CM = Combined Method) across three models. Gemma and Qwen results are mean over five runs; GPT results are based on one run. A dash (‘–’) indicates that the metric is undefined in at least one of the runs due to division by zero.
@@
Prompt Used for Vulnerability Detection
The prompt shown here is used when running LFM as well as for evaluating the vulnerability detection capabilities of the LLM for the experiments reported in Section V. To instruct the model clearly, we provided a concise and explicit system instruction, guiding the model to behave strictly as a security expert and to output responses in a standardized format. The exact system-level instruction used to prime the LLM is defined as follows:
@@ Results
We present a detailed analysis of the experimental results from our evaluation of three LLMs from the Gemma, GPT, and Qwen families across the four datasets (Results are presented separately for SVENC and SVENP). All findings are summarized in Table I, which reports the mean performance metrics over five independent runs for the Gemma and Qwen models using five different random seeds. The standard deviations between runs are minimal, so they are not included in the table because of space constraints. For the GPT model, we report the results from a single run. Our analysis is structured on a per-dataset basis to highlight the varying effectiveness of each few-shot selection approach under different data distributions and model capabilities.
@@
Finally, one interesting insight was hard examples (LFM) in context try to bias the model towards greater recall and lower precision. As CM3 applies LFM on LFNN examples keeping first LFM’s result as context, it also keeps that trend mostly.
Different Variants of Learn-from-Mistakes (LFM)
TABLE II: Performance comparison of different LFM variations. Each cell shows the mean and standard deviation over five runs. For the LFM variation names, the following abbreviations are used: S = stacked, U = unstacked, 1× = one iteration, m× = more than one iteration, inc. = incorrect, corr. = correct, and gray = gray cases. ‘–’ indicates metrics that are undefined due to division by zero. Dataset LFM Variations Gemma-3-4b-it Qwen-2.5-Coder Accuracy Precision Recall F1 Score Accuracy Precision Recall F1 Score DiverseVul S-1× (inc.) 0.525 (±0.003) 0.840 (±0.056) 0.061 (±0.005) 0.114 (±0.009) 0.500 (±0.000) 0.500 (±0.000) 1.000 (±0.000) 0.667 (±0.000) U-1× (inc.) 0.502 (±0.013) 0.501 (±0.007) 0.940 (±0.025) 0.654 (±0.010) 0.502 (±0.016) 0.501 (±0.010) 0.844 (±0.063) 0.628 (±0.022) U-m× (inc.) 0.507 (±0.010) 0.504 (±0.005) 0.995 (±0.008) 0.669 (±0.003) 0.527 (±0.039) 0.516 (±0.025) 0.976 (±0.045) 0.674 (±0.009) U-m× (corr.) 0.597 (±0.053) 0.778 (±0.056) 0.295 (±0.189) 0.383 (±0.204) 0.621 (±0.065) 0.752 (±0.027) 0.365 (±0.199) 0.455 (±0.199) U-m× (gray) 0.513 (±0.000) 0.667 (±0.000) 0.053 (±0.000) 0.099 (±0.000) 0.598 (±0.061) 0.713 (±0.059) 0.341 (±0.226) 0.415 (±0.199) NodeMedic S-1× (inc.) 0.725 (±0.000) 0.725 (±0.000) 1.000 (±0.000) 0.840 (±0.000) 0.720 (±0.000) 0.723 (±0.000) 0.993 (±0.000) 0.837 (±0.000) U-1× (inc.) 0.273 (±0.004) 0.500 (±0.707) 0.001 (±0.003) 0.003 (±0.006) 0.274 (±0.005) 0.333 (±0.422) 0.004 (±0.006) 0.009 (±0.011) U-m× (inc.) 0.277 (±0.000) - 0.000 (±0.000) 0.000 (±0.000) 0.272 (±0.003) 0.000 (±0.000) 0.000 (±0.000) U-m× (corr.) 0.725 (±0.000) 0.725 (±0.000) 1.000 (±0.000) 0.840 (±0.000) 0.724 (±0.002) 0.727 (±0.001) 0.991 (±0.003) 0.839 (±0.001) U-m× (gray) 0.506 (±0.003) 0.765 (±0.005) 0.460 (±0.000) 0.574 (±0.001) 0.467 (±0.100) 0.748 (±0.022) 0.412 (±0.245) 0.493 (±0.158) PrimeVul S-1× (inc.) 0.516 (±0.013) 0.726 (±0.186) 0.437 (±0.459) 0.336 (±0.270) 0.500 (±0.000) 0.500 (±0.000) 1.000 (±0.000) 0.667 (±0.000) U-1× (inc.) 0.507 (±0.010) 0.529 (±0.054) 0.773 (±0.352) 0.551 (±0.212) 0.497 (±0.027) 0.489 (±0.042) 0.466 (±0.137) 0.470 (±0.092) U-m× (inc.) 0.496 (±0.009) 0.397 (±0.221) 0.034 (±0.022) 0.061 (±0.038) 0.501 (±0.005) 0.502 (±0.005) 0.864 (±0.219) 0.621 (±0.077) U-m× (corr.) 0.499 (±0.002) 0.499 (±0.001) 0.886 (±0.223) 0.626 (±0.079) 0.489 (±0.010) 0.386 (±0.090) 0.034 (±0.010) 0.062 (±0.018) U-m× (gray) 0.499 (±0.002) 0.496 (±0.009) 0.366 (±0.284) 0.357 (±0.202) 0.498 (±0.017) 0.529 (±0.141) 0.246 (±0.241) 0.267 (±0.198) SVENC S-1× (inc.) 0.500 (±0.000) 0.500 (±0.000) 1.000 (±0.000) 0.667 (±0.000) 0.500 (±0.000) 0.500 (±0.000) 1.000 (±0.000) 0.667 (±0.000) U-1× (inc.) 0.500 (±0.000) - 0.000 (±0.000) 0.000 (±0.000) 0.500 (±0.000) 0.000 (±0.000) 0.000 (±0.000) U-m× (inc.) 0.500 (±0.000) - 0.000 (±0.000) 0.000 (±0.000) 0.500 (±0.000) 0.000 (±0.000) 0.000 (±0.000) U-m× (corr.) 0.504 (±0.009) 0.503 (±0.006) 0.960 (±0.069) 0.659 (±0.013) 0.493 (±0.009) 0.496 (±0.005) 0.911 (±0.056) 0.642 (±0.016) U-m× (gray) 0.478 (±0.000) 0.474 (±0.000) 0.400 (±0.000) 0.434 (±0.000) 0.502 (±0.011) 0.080 (±0.109) 0.114 (±0.142) SVENP S-1× (inc.) 0.500 (±0.000) 0.500 (±0.000) 0.905 (±0.013) 0.644 (±0.003) 0.500 (±0.000) 0.500 (±0.000) 1.000 (±0.000) 0.667 (±0.000) U-1× (inc.) 0.558 (±0.067) 0.571 (±0.083) 0.600 (±0.127) 0.573 (±0.050) 0.590 (±0.051) 0.607 (±0.074) 0.547 (±0.111) 0.567 (±0.062) U-m× (inc.) 0.540 (±0.033) 0.567 (±0.079) 0.447 (±0.130) 0.483 (±0.074) 0.571 (±0.054) 0.569 (±0.065) 0.721 (±0.132) 0.624 (±0.041) U-m× (corr.) 0.524 (±0.023) 0.522 (±0.018) 0.621 (±0.154) 0.557 (±0.063) 0.629 (±0.038) 0.712 (±0.045) 0.437 (±0.133) 0.528 (±0.110) U-m× (gray) 0.547 (±0.062) 0.549 (±0.064) 0.742 (±0.145) 0.619 (±0.034) 0.626 (±0.038) 0.767 (±0.120) 0.437 (±0.175) 0.519 (±0.106)
@@
Table II shows a comprehensive breakdown of performance across different LFM configurations. It is evident that the method’s effectiveness is highly sensitive to its parameterization and the specific characteristics of the dataset.